If Eratosthehne were alive in the U.S. today, he could repeat his calculation of the circumference of Earth. Wichita, Kansas, is 345 miles due north of Fort Worth, Texas. At a given moment, the sun is 5 degrees lower in the sky in Wichita than Fort Wroth. From this information, what estimate would Eratosthenes obtain for the Earth's circumference?
Answer by stanbon(75887)

5/345=x/180
x=12420 miles (length pf perimeter from pole to pole)
Circumference = 2(12480)=24960 miles
Cheers,
Stan H.